# Which Similarity-Sensitive Entropy?

### v1

@@ 4.2 Trends with LCR and VS
Fig. 5 presents comparisons among LCR, VS, and the number of HDBSCAN clusters across all datasets, with imaging and tabular datasets evaluated separately. As observed for MNIST, both LCR and VS were generally larger than the cluster count. The principal exceptions were the CIFAR image datasets, where the absence of meaningful substructure produced hundreds of very small clusters. Correlations between the number of clusters and LCR, and between the number of clusters and VS, were both low ($R^{2}=0.05$ and $0.06$, respectively), indicating that these entropy measures capture information not captured by clustering. In contrast, LCR and VS correlated strongly with each other for both imaging and tabular datasets ($R^{2}=1.00$ and $0.90$, respectively). Fig. 7 shows results for $q=0$ and $\infty$.
Figure 5: Correlations among LCR, VS, and HDBSCAN for (a) imaging and (b) tabular datasets. Each point represents a dataset. All entropies are effective-number forms at $q=1$. Red line = linear regression fit; gray line = 1:1 (off the scale to the left for (a), far right). Select datasets are labeled. (c) UMAPs for labeled datasets. Gray points = unclustered elements.
@@
When $k$ is extremely small, all $z_{ij}$ approach 1, collapsing the system to a single individual for both LCR and VS (Fig. 6a), yielding a perfect correlation ($R^{2}=1.00$). When $k$ is extremely large, all off-diagonal $z_{ij}$ approach 0, making every element essentially unique; consequently, both LCR and VS converge to the number of elements in the dataset (Fig. 6a), and the correlation again is perfect ($R^{2}=1.00$) (Fig. 6c). Only at intermediate $k$ values do LCR and VS provide independent information (Fig. 6b).
@@
Figure 6: LCR vs. VS at $q=1$ at a very small half distance, the half distance with the minimum $R^{2}$, and a very large half distance, together with the $R^{2}$, covariance, and correlation for $q=0,1,$ and $\infty$ for (a-b) imaging and (c-d) tabular datasets. (a) and (c): red line = linear regression fit; gray line = 1:1. Each point represents a dataset. (b) and (d): gray lines = $k$s for (a) and (c).
@@ 5 Discussion
Quantification is important across most if not all complex systems. This includes ML datasets, where the fact that most elements are generally unique—no repeated images (in image datasets) or rows (in tabular datasets)—renders traditional entropy (Shannon entropy and the other Rényi entropies) uninformative. This is because traditional entropy measures information encoded by the shape of the elements’ frequency distribution, but when every element is unique this distribution is flat. Traditional entropy is used to measure the size distribution of classes (class balance), but there is no natural way to use it to measure the dataset based on elements themselves.
5.1 Richer dataset descriptions
@@
A popular option is to use clustering algorithms, to quantify substructure as the number of clusters (and possibly their relative sizes, which can be measured by traditional entropy). Like LCR and VS, which require a choice of similarity measure and $q$, clustering algorithms also require choices. For example, HDBSCAN, which we used here, requires choosing min_cluster_size, min_samples, etc. Conceptually, the main difference is that similarity-sensitive entropy accounts for similarities both within and between clusters, whereas most clustering algorithms lack this ability: effectively, the within-cluster similarity is always 1 and the between-cluster similarity is always 0.
@@
The exception is fuzzy clustering, in which each element can belong to more than one cluster, with each element being parameterized by membership strength for one or more clusters. One can view LCR as carrying fuzzy-clustering to its logical limit, with each unique element being its own cluster and the similarities giving other elements’ strength of membership to that cluster, and thereby contributing to the size of that cluster; this is a fuzzy-clustering interpretation of the ordinariness, $Z\mathbf{p}_{i}$ (a big cluster that many unique elements belong to fairly strongly is “ordinary”) (Fig. 2c). This is another way of seeing that LCR is the entropy of a system after accounting/adjusting/correcting for the similarity among its elements. Traditional entropy accounts for frequency; LCR accounts for both frequency and similarity. In this way, LCR is a richer, more accurate measure of the information in a system.
@@
Regardless of whether they are represented as effective numbers (as we [nguyen2023greylock] and others [hill1973diversity, jost2006entropy, leinsterEntropyDiversityAxiomatic2020] have advocated) or as bits of information, how should LCR values be interpreted? LCR is always bounded above by the corresponding traditional entropy [leinsterEntropyDiversityAxiomatic2020]; for example, at $q=1$ LCR never exceeds the Shannon entropy. Similarity saps individual elements of some of their uniqueness.
@@
How do LCR and VS compare with the number of clusters identified by HDBSCAN? Predicting the relationship is difficult: large clusters whose peripheral elements are relatively distinct tend to increase the LCR/VS values, whereas strong similarities among elements—both within a cluster and across neighboring clusters—tend to reduce them (see Fig. 1). Indeed, the many datasets examined exhibited a pronounced difference in magnitude between tabular and imaging datasets when comparing the HDBSCAN cluster count with LCR and VS. For tabular data sets, both LCR and VS almost always surpassed the HDBSCAN cluster count, often by one to two orders of magnitude, reflecting that many rows remain unclustered by hard clustering.
@@
In contrast, for imaging datasets, LCR and VS were typically one to two orders of magnitude lower than the number of clusters, indicating substantial similarity among clusters that hard clustering does not capture (Fig. 2c). We interpret these systematic differences as evidence of the advantages of similarity‑sensitive entropy relative to hard clustering, for describing a system’s structure. We do not regard this as a failure of HDBSCAN; rather, we note as a limitation of our study that we did not explore a range of HDBSCAN hyperparameters, since the goal was simply to illustrate the general relationship between similarity‑sensitive entropy and clustering. However, we note that by the “eye test,” UMAP plots validate HDBSCAN’s clustering, suggesting hyperparameter search would be unlikely to meaningfully affect our findings.
5.2 The effect of scaling similarity
@@
We demonstrated that the relationship between LCR and VS depends strongly on how the similarity‑matrix values are scaled, which we illustrated using the half‑distance parameter $k$. Recall that similarity between two elements ranges from 0 to 1; scaling modifies this relationship by raising the values to a power, thereby shifting the curve upward or downward.
@@
Why should one need to scale the similarity between elements in the first place? Should scaling be permitted at all? When similarity is derived from external knowledge or a principled model, we argue that scaling should not be applied: the similarity function is defined a priori, its output for any given pair is fixed, and the similarity matrix simply collates these values for all pairs. An example is binding similarity in immune repertoires, defined as the relative dissociation constants of two antibodies (or TCRs) for a given antigen (or set of antigens) [aroraRepertoirescaleMeasuresAntigen2022]. This is the unique definition that preserves biophysical additivity; arbitrary scaling would break that relationship and thereby produce an objectively incorrect similarity matrix.
@@
We note that the parameter $q$ offers a loosely analogous scaling to that used in traditional (and similarity-sensitive) entropy, allowing investigators to adjust the emphasis on frequency. To avoid the impression of arbitrariness, remember that specific integer values of $q$ correspond to well‑known statistics; for example, $q=2$ corresponds to Simpson’s index, which reflects the statistics of collision frequency. It is possible that alternative values of $k$ could similarly be linked to other quantities of interest; this is left for future work.
@@
A key feature of scaling is that it interpolates similarity‑sensitive entropy between 1 (for small $k$) and traditional entropy (for large $k$). When $k$ is small, the off‑diagonal entries of the similarity matrix approach 1, implying that all unique elements are mutually similar; in the limit the system behaves as if it contained a single unique element. Conversely, for large $k$ the off‑diagonal entries approach zero, yielding the identity matrix in the limit; the unique elements become completely dissimilar, and traditional entropy is recovered. In a system where the $n$ elements are all unique—typically the case for ML data sets, including the imaging and tabular data sets examined here—the effective‑number form of traditional entropy equals the dataset size. Consequently, for such uniform systems $k$ interpolates both LCR and VS from 1 to $n$. Other scaling schemes may exhibit the same property.
@@
So: for heuristic similarity functions, which $k$ should be selected? Qualitatively, the answer is a value roughly midway between the extremes, but our quantitative findings indicate that the optimal choice is not obvious. We found that the correlation between LCR and VS is minimized at $k=0.2$ for tabular data sets and $3.4$ for imaging data sets—not at any reasonable definition of the midpoint of the range (defined for example as the geometric mean of the $k$ values where the correlation first falls to 0.99 on each side). Moreover, the shapes of these curves, with their multiple extrema, do not appear to follow a simple parametric form, making it difficult to predict precisely where the two metrics are maximally mutually informative. We also observed that even at this $k$, the $R^{2}$ between LCR and VS remained relatively high (0.680.68)—perhaps to be expected, since both metrics quantify the effective number of elements using the same underlying similarity measure, albeit in different ways (see Section 6.1). The dependence on $k$, however, was quite strong.
@@
Our results suggest first testing whether the entropy lies near 1 or $n$; if so, to explore different $k$ values to delineate a non‑trivial range; and finally to select a $k$ approximately in the middle of this range (on a logarithmic scale). Both the choice of similarity measure and the value of $k$ have effects on the resulting entropy that cannot be ignored. Thus to some extent, the information in a system is in the eye of the beholder.
@@ 6 Conclusion
6.1 LCR vs. VS
Our results support defaulting to LCR when the goal is to measure the similarity‑adjusted entropy of a system, and to use VS only in one of two special cases: (i) when the elements are usefully thought of as linear combinations of a set of mutually orthogonal ur‑elements/eigen‑elements, or (ii) when the system possesses a quantum‑mechanical character (see below). If the aim is a more general multidimensional characterization, there is no clear reason to avoid employing both metrics; keeping in mind that they are likely to correlate, our findings recommend selecting $k$ within the non‑trivial range (see Section 5.2) to maximize their independence and thereby their combined informativeness.
@@
From a practical perspective, LCR has three advantages. First, LCR is less computationally expensive whenever elements appear multiple times (i.e., when $\mathbf{p}$ is not flat/uniform), because the similarity matrix for LCR has order equal to the number of unique elements, whereas VS requires a matrix whose order equals the total number of elements; both calculations scale as the square of the matrix order. Empirically, we observed a speed advantage of greylock over vendi-score even when all elements were unique, a situation that frequently occurs in ML datasets (Fig. 8). This performance gap may narrow or widen as algorithmic improvements are introduced.
@@
Figure 8: Compute times in seconds on all 53 datasets for LCR (via greylock) vs. VS (via vendi-score). Imaging datasets were all 10,000 images.
@@
Second, LCR does not require the similarity matrix to be PSD, whereas VS does. Similarity matrices need not be PSD; a perfectly reasonable set of similarity measurements can produce a matrix that fails the PSD condition, which would complicate the use of VS. Even symmetric similarity matrices—the most common type, satisfying $z_{ij}=z_{ji}$, so that similarity is reciprocal—are not guaranteed to be PSD, even when their entries obey the triangle inequality. The prevalence of such non‑PSD cases in practice is unknown; we note that all similarity matrices examined for our datasets were PSD. Thus, PSD‑ness is irrelevant for LCR but must be considered when applying VS.
@@
And third, LCR in its various forms [chao2010phylogenetic, chaoAttributediversityApproachFunctional2019] is part of an extensively characterized framework [leinsterEntropyDiversityAxiomatic2020] that provides similarity‑sensitive analogs of mutual information, relative entropy, and cross‑entropy and offers methods for partitioning the overall similarity‑sensitive entropy ($\gamma$) into within‑class ($\alpha$) and between‑class ($\beta$) components. In entropic terms, $\alpha$ is the (effective‑number form of the) joint entropy of the elements together with their class label; $\beta$ is the mutual information between the elements and the class label; and $\gamma$ is the marginal entropy of the elements irrespective of class. In this work we computed LCR in its $\gamma$ form; when only a single class is present, $\gamma$ coincides with $\alpha$. Measuring the full $\alpha$, $\beta$, and $\gamma$ decomposition across ML classes is a promising direction for future research. (The greylock implementation can compute $\alpha$, $\beta$, and $\gamma$.)
@@
The necessary mathematics for extending this decomposition to VS exists, because VS is based on von Neumann (quantum) entropy, for which quantum mutual information, quantum relative entropy, and related quantum quantities are well understood [nielsen2010quantum]. However, computing these quantities requires a matrix that simultaneously encodes each element and its class membership. If there are $n$ total elements and $k$ classes, the VS decomposition demands either an $nk\times nk$ matrix or an $n\times k\times n\times k$ tensor (which can be reduced to $n\times n$ and $k\times k$ matrices by taking a partial trace). In contrast, LCR requires only the similarity matrix $Z$. Consequently, VS needs an explicit notion of similarity both between classes and between elements, whereas in LCR any class‑level similarity arises purely from the element‑level similarities within the classes. In principle, the VS formulation permits two classes that contain exactly the same individuals to nonetheless have zero similarity—a scenario that cannot occur under LCR, in which the similarity between classes is purely a function of similarities among the elements they contain.
@@
In sum, LCR and VS provide complementary descriptions of similarity‑sensitive entropy, a unifying concept that is likely to find applications in many domains where traditional entropy is a mainstay. We anticipate that the deeper understanding of their relationship and properties as presented here will aid such investigations.

### v2

@@ 4.2 Trends with LCR and VS
Fig. 5 presents comparisons among LCR, VS, and the number of HDBSCAN clusters across all datasets (for default $k$ at $q=1$; Figs. 6 and 7 investigate other $k$s and $q$s), with imaging and tabular datasets evaluated separately. As observed for MNIST, both LCR and VS were generally larger than the cluster count. The principal exceptions were the CIFAR image datasets, where the absence of meaningful substructure produced hundreds of very small clusters. Correlations between the number of clusters and LCR, and between the number of clusters and VS, were both low ($R^{2}=0.05$ and $0.06$, respectively), indicating that these entropy measures capture information not captured by clustering. In contrast, LCR and VS correlated strongly with each other for both imaging and tabular datasets ($R^{2}=1.00$ and $0.90$, respectively). Fig. 7 shows results for $q=0$ and $\infty$.
Figure 5: Correlations among LCR, VS, and HDBSCAN for (a) imaging and (b) tabular datasets. Each point represents a dataset. All entropies are effective-number forms at $q=1$. Red line = linear regression fit; gray line = 1:1 (off the scale to the left for (a), far right). Select datasets are labeled. (c) UMAPs for labeled imaging datasets. Gray points = unclustered elements.
@@
When $k$ is very small, all $z_{ij}$ approach 1, collapsing the system to a single individual for both LCR and VS (Fig. 6a and c, left), yielding a perfect correlation ($R^{2}=1.00$). When $k$ is very large, all off-diagonal $z_{ij}$ approach 0, making every element essentially unique; consequently, both LCR and VS converge to the number of elements in the dataset, and the correlation again is perfect ($R^{2}=1.00$) (Fig. 6a and c, right). Only at intermediate $k$ values do LCR and VS provide independent information (Fig. 6a and c, middle).
@@
Figure 6: LCR vs. VS at $q=1$ at a very small half distance, an intermediate half-distance for which the correlation is low, and a very large half distance, together with the $R^{2}$, covariance, and correlation for $q=0,1,$ and $\infty$ for (a-b) imaging and (c-d) tabular datasets. For tabular datasets, the intermediate half-distance is the one simultaneously minimizing the $R^{2}$, the correlation, and maximizing the covariance at $q=1$. For image datasets, the intermediate half-distance does not coincide with the global minimum of the correlation at $q=1$, because VSs and LCRs already come very close to 10,000 at that global minimum, so we chose a smaller value. (a) and (c): red line = linear regression fit; gray line = 1:1. Each point represents a dataset. (b) and (d): vertical gray lines = $k$s for (a) and (c).
@@ 5 Discussion
Quantifying the information contained in a system is essential for virtually any complex system. This includes ML datasets, where the fact that most elements are generally unique—no repeated images (in image datasets) or rows (in tabular datasets)—renders traditional entropy (Shannon entropy and the other Rényi entropies) uninformative. This is because traditional entropy measures information encoded by the shape of the elements’ frequency distribution, but when the elements each appear only once, this distribution is flat (Fig. 1): there is no additional information there. Traditional entropy is used to measure the distribution of class sizes, which often vary; this is class balance. However, there is no obvious way to use class balance to measure the the information contained within or between classes, which is a property of the elements themselves.
5.1 Richer measures of the information that ML datasets contain
@@
One way to get at this missing information is to use clustering algorithms, to quantify dataset substructure as the number of clusters and possibly clusters’ relative sizes (which can be measured by traditional entropy). Like LCR and VS, which require a choice of similarity measure and $q$, clustering algorithms require choices. For example, HDBSCAN requires choosing min_cluster_size, min_samples, etc. Conceptually, the main difference between clustering and S-entropy is that S-entropy accounts for similarities both within and between clusters, whereas most clustering algorithms (and classifiers) lack this ability: effectively, the within-cluster (or intra-class) similarity is always 1 and the between-cluster similarity is always 0.
@@
The exception is fuzzy clustering, in which each element can belong to more than one cluster, with each element being parameterized by its strength of membership for one or more clusters. One can view LCR as carrying fuzzy-clustering to its logical limit, with each unique element being its own cluster and the similarities giving other elements’ strength of membership to that cluster and thereby contributing to the size of that cluster. This is a fuzzy-clustering interpretation of the ordinariness, $Z\mathbf{p}_{i}$: a big cluster that many unique elements belong to fairly strongly is “ordinary” (Fig. 2c). This perspective provides another way of seeing that LCR is the entropy of a system after accounting/adjusting/correcting for the similarity among its elements.
@@
Traditional entropy accounts for frequency; LCR accounts for both frequency and similarity. In this way, LCR is a richer, more accurate measure of the information in a system. Note that, regardless of whether they are represented as effective numbers (as we [nguyen2023greylock] and others [hill1973diversity, jost2006entropy, leinsterEntropyDiversityAxiomatic2020] have advocated) or as bits of information, LCR is always bounded above by the corresponding traditional entropy [leinsterEntropyDiversityAxiomatic2020]; for example, at $q=1$ LCR never exceeds Shannon entropy. Similarity saps individual elements of some of their uniqueness, but in doing so, it reveals differences between datasets that traditional entropy cannot (Fig. 1).
@@
How do LCR and VS compare with the number of clusters identified by HDBSCAN? Predicting the relationship is difficult: large clusters whose peripheral elements are relatively distinct tend to increase the LCR/VS values, whereas strong similarities among elements—both within a cluster and across neighboring clusters—tend to reduce them (see Figs. 1, 5c). Indeed, many datasets examined exhibited a pronounced difference in magnitude between tabular and imaging datasets when comparing the HDBSCAN cluster count with LCR and VS. For tabular data sets, both LCR and VS almost always surpassed the HDBSCAN cluster count, often by one to two orders of magnitude, reflecting that many rows remain unclustered by hard clustering.
@@
In contrast, for imaging datasets, LCR and VS were typically one to two orders of magnitude lower than the number of clusters, indicating substantial similarity among clusters that hard clustering does not capture (Fig. 2c). We interpret these systematic differences as evidence of the advantages of similarity‑sensitive entropy relative to hard clustering, for describing a system’s structure. We do not regard this as a failure of HDBSCAN; we note as a limitation of our study that we did not explore a range of HDBSCAN hyperparameters, since the goal was simply to illustrate the general relationship between similarity‑sensitive entropy and clustering. However, we note that by the “eye test,” UMAP plots validate HDBSCAN’s clustering, suggesting hyperparameter search would be unlikely to meaningfully affect our findings.
5.2 The effect of scaling the similarity measure/similarity matrix
@@
We demonstrated that the relationship between LCR and VS depends strongly on how the similarity‑ matrix values are scaled, which we illustrated by introducing the half‑distance parameter $k$. Recall that similarity between two elements ranges from 0 to 1; scaling modifies this relationship by raising the values to a power, thereby shifting the curve upward or downward in the middle, while maintaining the anchoring at each end (at 0 and 1).
@@
Why should one need to scale the similarity between elements in the first place? Should scaling be permitted at all? When similarity is derived from external knowledge or some principle or outside requirement, we argue that scaling should not be applied: the similarity function is defined a priori, its output for any given pair is fixed, and the similarity matrix simply collects these values for all pairs. An example is binding similarity in immune repertoires, defined as the relative dissociation constants of two antibodies (or TCRs) for a given antigen (or set of antigens) [aroraRepertoirescaleMeasuresAntigen2022]. This is the unique definition that preserves biophysical additivity; arbitrary scaling would break this relationship and thereby produce a similarity matrix that is objectively incorrect.
@@
Such choices arise often in measurement. For example, we note that the parameter $q$ offers a loosely analogous scaling to that used in traditional (and similarity-sensitive) entropy, allowing investigators to adjust the emphasis on frequency. Lest choices feel too arbitrary, remember that specific integer values of $q$ correspond to well‑known statistics; for example, $q=2$ corresponds to Simpson’s index, which reflects the statistics of collision frequency. It is possible that alternative values of $k$ could similarly be linked to other quantities of interest. Investigating this possibility is left for future work.
@@
A key feature of scaling is that it interpolates similarity‑sensitive entropy between 1 (for small $k$) and traditional entropy (for large $k$). When $k$ is small, the off‑diagonal entries of the similarity matrix approach 1, implying that all unique elements are mutually similar; in the limit the system behaves as if it contained a single unique element. Conversely, for large $k$ the off‑diagonal entries approach zero, yielding the identity matrix in the limit; the unique elements become completely dissimilar, and traditional entropy is recovered. In a system where the $n$ elements are all unique—typically the case for ML data sets, including the imaging and tabular data sets examined here—the effective‑number form of traditional entropy equals the dataset size. Consequently, for such uniform systems $k$ interpolates both LCR and VS from 1 to $n$. Other scaling approaches may exhibit the same property.
@@
So: for heuristic similarity functions, which $k$ should one choose? Qualitatively, the answer is a value roughly midway between the extremes (Fig. 6). However, our quantitative findings indicate that the optimal choice is not obvious. We found that the correlation between LCR and VS is minimized at $k=0.2$ for tabular data sets and $3.4$ for imaging data sets—not at any obvious definition of the midpoint of the range (defined for example as the geometric mean of the $k$ values where the correlation first falls to 0.99 on each side). Moreover, the shapes of these curves, with their multiple extrema, do not appear to follow a simple parametric form, making it difficult to predict precisely where the two metrics are maximally mutually informative. We also observed that even at the $k$ that minimizes the LCR-VS correlation, the $R^{2}$ between LCR and VS remained relatively high in absolute terms, at $0.68$. Perhaps this was to be expected, since both metrics quantify the effective number of elements using the same underlying similarity measure, albeit in different ways (see Section 6). The dependence on $k$, however, was quite strong.
@@
Our results suggest first testing whether the entropy lies near 1 or $n$; if so, to explore, if possible, different $k$ values to delineate a non‑trivial range; and finally to select a $k$ approximately in the middle of this range (on a logarithmic scale). Both the choice of similarity measure and the value of $k$ have effects on the resulting entropy that cannot be ignored. Thus to some extent, the information in a system is in the eye of the beholder.
@@ 6 Conclusion
So: LCR or VS? Our results support defaulting to LCR when the goal is to measure the similarity‑adjusted entropy of a system, and to use VS only in one of two special cases: (i) when the elements are usefully thought of as linear combinations of a set of mutually orthogonal ur‑elements/eigen‑elements, or (ii) when the system possesses a quantum‑mechanical character (see below). If the aim is a more general multidimensional characterization, there is no clear reason to avoid employing both metrics; keeping in mind that they are likely to correlate, our findings recommend selecting $k$ within the non‑trivial range (see Section 5.2) to maximize their independence and thereby their combined informativeness.
@@
From a practical perspective, LCR has three advantages. First, it is less computationally expensive whenever elements appear multiple times (i.e., when $\mathbf{p}$ is not flat/uniform), because the similarity matrix for LCR has order equal to the number of unique elements, whereas VS requires a matrix whose order equals the total number of elements; both calculations scale as the square of the matrix order. Empirically, we observed a speed advantage of greylock over vendi-score even when all elements were unique, a situation that frequently occurs in ML datasets (Fig. 8). This performance gap may narrow or widen as codebases evolve.
@@
Figure 8: Compute times for all 53 datasets for LCR (using the greylock Python package [nguyen2023greylock]) vs. VS (using vendi-score). Imaging datasets were all 10,000 images. Smaller datasets were faster. Note the different timescales.
@@
Second, LCR does not require the similarity matrix to be PSD, whereas VS does. Similarity matrices need not be PSD; a perfectly reasonable set of similarity measurements can produce a matrix that fails the PSD condition, which would complicate the use of VS. Even symmetric similarity matrices—the most common and most natural type, satisfying $z_{ij}=z_{ji}$, so that similarity is reciprocal—are not guaranteed to be PSD, even when their entries obey the triangle inequality. The prevalence of such non‑PSD cases in practice is unknown; we note that all similarity matrices examined for our datasets were PSD (a consequence of how we chose our similarity measures; this follows from Theorem 2.11 in [leinster2013magnitude]). Thus, PSD‑ness is irrelevant for LCR but must be checked to apply VS.
@@
And third, LCR in its various forms [chao2010phylogenetic, chaoAttributediversityApproachFunctional2019] (Table 1) is part of an extensively characterized framework [leinsterEntropyDiversityAxiomatic2020] that provides similarity‑sensitive analogs of mutual information, relative entropy, and cross‑entropy and offers methods for partitioning the overall similarity‑sensitive entropy ($\gamma$) into within‑class ($\alpha$) and between‑class ($\beta$) components. In entropic terms, $\alpha$ is the (effective‑number form of the) joint entropy of the elements together with their class label; $\beta$ is the mutual information between the elements and the class label; and $\gamma$ is the marginal entropy of the elements irrespective of class. In this work we computed LCR in its $\gamma$ form; when only a single class is present, $\gamma$ coincides with $\alpha$. Measuring the full $\alpha$, $\beta$, and $\gamma$ decomposition across ML classes is a promising direction for future research. (The greylock implementation computes $\alpha$, $\beta$, and $\gamma$.)
@@
The necessary mathematics for extending this decomposition to VS exists, because VS is based on von Neumann (quantum) entropy, for which quantum mutual information, quantum relative entropy, and related quantum quantities are well understood [nielsen2010quantum]. However, computing these quantities requires a matrix that simultaneously encodes each element and its class membership. If there are $n$ total elements and $k$ classes, the VS decomposition demands either an $nk\times nk$ matrix or an $n\times k\times n\times k$ 4-dimensional tensor (which can be reduced to $n\times n$ and $k\times k$ matrices by taking a partial trace). In contrast, LCR requires only the similarity matrix $Z$. Consequently, VS needs an explicit notion of similarity both between classes and between elements, whereas in LCR any class‑level similarity arises purely from element‑level similarities within the classes. In principle, the VS formulation would permit the confusing situation in which two classes contain exactly the same individuals but nonetheless have zero similarity—a scenario that cannot occur under LCR, in which the similarity between classes is purely a function of similarities among the elements they contain.
@@
In sum, LCR and VS provide complementary descriptions of similarity‑sensitive entropy, a unifying concept that is likely to find applications in many domains in which traditional entropy is currently a mainstay. We anticipate that the deeper understanding of their relationship and properties through work such as that presented here will aid such investigations.
7 Acknowledgements
This work was supported by the Gordon and Betty Moore Foundation and by the NIH under grants R01HL150394, R01HL150394-SI, R01AI148747, and R01AI148747-SI.